Please fill the table for your shift with this information:

  • PROGRAM DONE: please write the type of measurements you have done, with reference to run number range and to logbook pages
  • BEAM INFO: note periods with no beam, report related problems
  • ASIDE ACTIVITIES: please add information on aside activities such as debugging… put references to logbook pages
  • OBSERVED PROBLEMS: add Meaningful information of problems encountered in your shift

Remember to also send an email to the dream mailing list.

Shift Date 2/12/2014
Shift Time 00-08
Shifter Names Michele C, Alan
Program Done Cu module (aluminized) theta scan
Beam Info H8C.DREAM.001 20 GeV electrons(?)
Aside Activities Wrote a root macro (~/DreamDaq/macro_root/cu_dream_avg.cxx) to automate the fit of the two Cu_dream_tot_* distributions. TWiki cleanup. Added cron job to backup data files at TTU via globus every hour.
Observed Problems 4:10 pcdreamus froze, we rebooted it and tried to take a run (10123) that wen way too fast (20k ev in a few seconds). We rebooted pcdreamdaq2 and the problem disappeared (???)
7:10 beam lost far a few minutes

Shift Date 2/12/2014
Shift Time 08-16
Shifter Names Fangqi, John
Program Done Angle scan at 1 degree tilt
Beam Info H8C.DREAM.001 20 GeV electrons
Aside Activities  
Observed Problems The machine is down for 2 hours.

Shift Date 2/12/2014
Shift Time 16-24
Shifter Names Silvia, Evelin
Program Done Continue Angle scan at 1 degree tilt, without and with PS
Beam Info H8C.DREAM.001 20 GeV electrons
Aside Activities  
Observed Problems 22:00 beam lost

Shift Date 3/12/2014
Shift Time 00-08
Shifter Names laura, michele c
Program Done Continue Angle scan at 1 degree tilt, with PS
Beam Info H8C.DREAM.001 20 GeV electrons
Aside Activities Added labels to Pb matrix cables
Observed Problems 22:00 beam off
2.00 beam back but with low intensity (30 triggers/spill)
2.50 beam back with normal intensity (700 ev/spill) data taking restarted
6:00 beam off (Machine Development)

Shift Date 3/12/2014
Shift Time 16-24
Shifter Names Silvia, Richard
Program Done restart after MD, high energy wobbling, new pion runs, setup beam, VME crate crashed, problems to restart. Roberto and Fabrizio solved. runs with DRS, they set up the timing of DRS in order to take fast calorimeter signals and low (150ns late) neutron counter signals, due to longer cables. DRS sampling set at 2.5GHZ. Started calibration Pb matrix with electrons and readout with ADC
Beam Info 180 GeV pions, starting to setup electrons
Aside Activities  
Observed Problems V262 IO register crashed both the running and the spare module. Problems also with the disk (Alan is taking care about that)

Shift Date 6/12/2014
Shift Time 16-24
Shifter Names Silvia,Fabrizio
Program Done redone 180GeV pions with DRS, high rate, so we took also a run with 100k events, came back to ADCs to continue the run at different pion energies. 40 GeV pi energy, low stat (70 ev per spill).
Beam Info 180 GeV pions, 40 GeV pions
Aside Activities  
Observed Problems  

-- MicheleCascella - 2014-11-25

Edit | Attach | Watch | Print version | History: r11 < r10 < r9 < r8 < r7 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r11 - 2014-12-06 - SilviaFranchino
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    DREAM All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback